OpenStudy (anonymous):

is -x+y-3=0 and 6x+y-1=0 is it parallel, coincident or intercecting?plese help

OpenStudy (tkhunny):

Calculate the Slope of each linear equation.

OpenStudy (tkhunny):

This is normally done by solving fir y and putting the equation in Slope-Intercept form. y = mx + b

OpenStudy (anonymous):

y=x+3 and y=-6x+1?

OpenStudy (mathmale):

good work! If the slopes are the same, the 2 lines are either parallel or coincident (one lies right on top of the other one) If the slopes are different, the two lines intersect. what's the importance of that?